# thanks: These two authors contributed equallythanks: These two authors contributed equally

### v2

@@ Abstract
Introduction.—Quantum systems interacting with Markovian environments are ubiquitous in the physical sciences. A main tool for studying their dynamics is the determinsitic Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) master equation (ME) Gorini et al. (1976); Lindblad (1976). This specifies the time evolution of the density matrix $\rho(t)$ as the system experiences both coherent and incoherent processes, with the latter involving leakage of state information to the environment Breuer and Petruccione (1999); Haroche and Raimond (2006); Breuer and Petruccione (2002); Attal et al. (2006). Despite its generality and wide use, the GKSL ME does not fully describe the quantum dynamics when the environment includes measurement devices, which convert a portion of the leaked information to usable form. The temporal evolution conditioned on the measurement record $m$ defines a quantum trajectory, in the ideal case  [^1] The unraveling into pure states can be performed whenever the continuous evolution in the GKSL ME is governed by a commutator with a non-Hermitian Hamiltonian, and the initial state is pure. expressed as the pure state $\rho_{m}(t)=\ket{\psi_{m}(t)}\bra{\psi_{m}(t)}$ |. Averaging $\rho_{m}(t)$ over the measurement record solves the GKSL ME Accardi et al. (2013). The identification of the $\rho(t)$ with the ensemble average of $\rho_{m}(t)$ is an example of an unraveling of the ME into a stochastic equation for the pure state $\rho_{m}(t)$ N. Bohr and Slater (1924); Carmichael (1993); Gardiner et al. (1992); Dalibard et al. (1992); Omnès (1994); Blanchard and Jadczyk (1995); Korotkov (1999); Plenio and Knight (1998); Hegerfeldt and Sondermann (1996); Daley (2014). Evidently, different measurement schemes correspond to different unravelings and lead to different ensembles of quantum trajectories. Unravelings thus provide information that is not available from the corresponding ME.
@@
Quantities that are linear in the density matrix $\rho(t)$, such as averages of observables, are fully determined by the GKSL ME and, therefore, are independent of the choice of the unraveling dictated by a given measurement scheme. In this Letter, we develop nonlinear measures to differentiate unravelings, thus opening a way to access the physics beyond the ME. We demonstrate that evaluating an expectation value of a physical observable for a specified quantum trajectory $\rho_{m}(t)$, performing a nonlinear operation on the obtained result, and averaging the result over the measurement record, yields a quantity that allows for distinguishing different unravelings of the same GKSL ME. We focus on a paradigmatic open quantum system, the resonance fluorescence of a two-level atom, and consider unravelings corresponding to direct photodetection and to homodyne/heterodyne detection.
@@
The unravelings.—Electron shelving Wineland and Dehmelt (1975); Dehmelt (1981) paved the way to the first observations of quantum jumps Nagourney et al. (1986); Sauter et al. (1986); Bergquist et al. (1986), followed by several atomic Basché et al. (1995); Peil and Gabrielse (1999); Gleyzes et al. (2007) and solid-state physics experiments Neumann et al. (2010); Vijay et al. (2011); Minev et al. (2019). The theoretical description of these investigations dates back to early works Cook and Kimble (1985); Kimble et al. (1986); Erber and Putterman (1985); Javanainen (1986); Schenzle et al. (1982); Cohen-Tannoudji and Dalibard (1986); Cook (1988); Hegerfeldt and Plenio (1992, 1993); Hegerfeldt (2009) that stimulated the development of quantum trajectory theory Keys and Wehr (2020); Barchielli and Belavkin (1991); Brun (2002); Barchielli (2006); Belavkin (1990); Belavkin and Staszewski (1991); Ghirardi et al. (1990, 1986); Carmichael (1993). Two unravelings of the GKSL ME that play a fundamental role in the understanding of the quantum trajectories are: (i) Poisson unraveling, related to direct photodetection and the so called quantum Monte Carlo wave function approach Dalibard et al. (1992); Davies (1976); Srinivas and Davies (1981); Zoller et al. (1987); Dum et al. (1992); and (ii) Wiener-type unraveling (the quantum state diffusion model proposed by Gisin and Percival) Gisin and Percival (1992); Carmichael (1999), relating conditional quantum dynamics to a continuous Wiener process Durrett (2019). In the context of atomic physics experiments, the distinct unravelings correspond to different photodetection schemes Breuer and Petruccione (2002), see Fig. 1. The Poisson unraveling is relevant for the direct photodetection experiments, while the continuous Wiener process arises in homodyne and heterodyne photodection schemes Carmichael (1993); Wiseman and Milburn (1993).
@@
The disparities between the experimental setups is reflected in the different nature of quantum trajectories Carmichael (1999). The Wiener process yields a continuous evolution of the system state |$\ket{\psi(t)}$. In contrast, the acts of direct photodetection at times $t_{1}<t_{2}<\ldots<t_{n}$ collapse the conditional wavefunction. The resulting time evolution of |$\ket{\psi(t)}$ is discontinuous and the final state at $t>t_{n}$ depends, in general, on a particular sequence of emission times |$\ket{\psi(t)}=\ket{\psi_{t_{1},\ldots,t_{n}}(t)}$. By collecting photon counting records, the experimenter effectively determines the quantum trajectory of the atom. The entanglement between the electromagnetic field and the atom is the key ingredient that allows for the inference of the atom’s state based on the photodetection events Carmichael (1999); Nha and Carmichael (2004).
@@
Source Master Equation and linear averages.—Our starting point is the GKSL ME of resonance fluorescence, governing the unconditional [PS: it is by definition unconditional evolution of the reduced system density matrix $\rho$,
@@
The results depicted in Fig. 2 substantiate the pivotal influence of the environment when collecting records of a strongly driven two-state atom and taking a sum over a collection of them. The two principal unravelings are presented in their ability to produce an ostensibly disparate ${\rm Var}(\sigma_{z})$, while the corresponding average inversion remains unchanged. For the direct photodetection Mollow (1975); Cook (1980, 1981), corresponding to the Poisson-type unraveling of the ME, we obtain an exact expression for ${\rm Var}(\sigma_{z})$ based on the waiting-time distribution Carmichael et al. (1989). For $\gamma t\gg 1$, the asymptotic expression for the variance, including first-order terms in $\gamma/\Omega$ of different frequencies, reads SM
@@
Time evolution of QTAV is significantly altered, see the middle panels of Fig. 2, when one places a beam splitter and a local oscillator in the environment, and the fluorescent signal interferes with the latter before photodetection [Fig. 1(c)], corresponding to the heterodyne detection and exemplifying Wiener-type unraveling. The frequency doubling is also in evidence although the contrast in the oscillations is visibly suppressed. The light scattered by the two-level emitter is squeezed in the field quadrature that is in phase with the mean scattered field amplitude $\propto\langle\sigma_{-}(t)\rangle$ Walls and Zoller (1981); Mandel (1982). The bottom panel in each frame shows that the QTAV responds differently to the detection of the squeezed vs. the direction of the antisqueezed quadrature of the fluorescent field, i.e., along an axis perpendicular to the equator of the Bloch sphere where quantum fluctuations are redistributed among the quadratures.
@@
using the fact that $d\widetilde{N}d\widetilde{N}=dN$ with the rate of the Poisson process being $\gamma\left\langle\sigma_{+}\sigma_{-}\right\rangle_{t}dt$ and in the last line using $u^{j}$ as the vector corresponding to $\mathcal{L}^{\dagger}[X_{j}]$, $l$ as $\sigma_{+}\sigma_{-}$, and $v^{j}$ as $\sigma_{+}X_{j}\sigma_{-}$. This is an ordinary differential equation which however does not close since it requires higher order moments such as $\mathbb{E}x_{i}x_{j}x_{k}$. Again, using the Itô product rule we can calculate the equation for the higher order moments to obtain a system of ordinary differential equations which still do not close. We can repeat this procedure to arbitrarily high order but at some point we have to truncate. It can be shown that this truncation is linear in the moments, which allows us to solve the system using traditional methods of solving linear systems of ODEs. The Wiener case Barchielli and Belavkin (1991); Gisin and Percival (1992) can be similarly approximately solved by using the Heisenberg equation Keys (2022)
@@
Since the intensity correlation of the scattered light reflects a nonexclusive probability of photocounting coincidences, the limited detector efficiency can be counterbalanced by increasing the number of photon “clicks” $N$ in the course of a long experimental run. Indeed, for the setup pictured in Fig. 1(a), we have concluded that the signal-to-noise ratio for $g^{(2)}(\gamma\tau\gg 1)$ in a window about one inverse of the coherence time scales with $\sqrt{N}$ SM . This allows the determination of ${\rm Var}(\sigma_{z})$ from single realizations as low as $10^{-3}$, the order of magnitude Monte-Carlo simulations indicate for $\eta\lesssim 0.05$. This order of magnitude can be increased using high numerical aperture collection systems Bruno et al. (2019); Natarajan et al. (2012) and efficient single-photon detectors Chou et al. (2017).
@@
This ability allows for an experimentally oriented test of the objective quantum state assumption via an EPR steering inequality Jones et al. (2007); Wiseman and Gambetta (2012). It would be interesting to extend our considerations to non-Markovian evolution case Barchielli and Gregoratti (2012). Finally, our conclusions are reflected by recent investigations of quantum many-body systems studied in the context of quantum computing and quantum simulation Preskill (2018); Altman et al. (2021); Fraxanet et al. (2022). Quantum trajectories arising due to multiple measurements of the system’s state, when analyzed by relevant (nonlinear) statistical measures such as entanglement entropy, exhibit phase transitions Skinner et al. (2019); Li et al. (2018, 2019); Chan et al. (2019); Sierant and Turkeshi (2022); Turkeshi et al. (2020); Lunt et al. (2021); Sierant et al. (2022), that are not evident in the average state Cao et al. (2019); Piccitto et al. (2022); Kolodrubetz (2023); Turkeshi et al. (2021) unless specifically tuned feedback mechanisms are employed Sierant and Turkeshi (2023a); Buchhold et al. ; Iadecola et al. (2023); Piroli et al. (2023); O’Dea et al. ; Ravindranath et al. (2023a, b); Sierant and Turkeshi (2023b).
@@
Appendix A EPR steering for system-environment entanglement
@@
$S(\rho^{c,D},\rho^{c,H})\equiv\overline{f_{1}(\rho^{c,D})}+\overline{f_{2}(\rho^{c,H})}\leq 1,$ (SA.39)
@@
Figure SA.3: EPR steering $S=S(\rho^{c,D},\rho^{c,H})$ value from the inequality (SA.39) for: $Y=10$ and $\eta=1$ in (a), $Y=30$ and $\eta=1$ in (b), $Y=30$ and $\eta=0.8$ in (c), and $Y=30$ and $\eta=0.6$ in (d).
Figure SA.3 plots the value of $S$ for different values of $Y$ and $\eta$. As noted in Wiseman and Gambetta (2012), for $S>1$ the experiment would rule out all theories of objective atomic state reduction. For unit-efficiency direct detection ($\eta=1$), the steady-state value of $S$ always exceeds unity. We note the envelope is primarily determined by the detection efficiency while the carrier frequency is set by $Y$. Figures SA.3(c, d) show that the envelope drops below unity for $\eta\lesssim 0.8$, satisfying the inequality (SA.39), whence making the experiment “not provably measurement dependent”. In contrast, we note that the feedback scheme adopted for direct photodetection in Wiseman and Gambetta (2012) guarantees $S>1$ for any value of $\eta$ and a perfect heterodyne detector.
A.1 Experimental setup and considerations
@@
We consider real experimental data based on the setup of Bruno et al. (2019) in order to highlight the different imperfections. The experimental setup consists of a Maltese-cross coupling of single neutral atoms of 87Rb. The atom is illuminated by a laser field with Rabi frequency $\Omega$ and detuning $\Delta$. The photons are collected by high-numerical aperture lenses, as shown in the main text, and the photon arrival times are measured [Fig. SA.4(a)] using single-photon avalanche photodiodes (APDs). From these measurements, the normalized second-order correlation function between the lenses L1 and L2 is computed [Fig. SA.4(b)]. The average atomic scattering rate on a single detector is $R_{\text{sca,det}}=9000$ counts/s.
Figure SA.4: (a) The arrival times of the photons coming from the lenses L1 and L2 are measured by avalanche photodiodes (APDs). (b) Normalized second-order correlation function for integration times of 65 (blue) or 1700 (black) seconds and fit using Eq. (SA.40) from a two-level system approximation including experimental imperfections (red). The gray shaded box corresponds to the time range where the signal-to noise of $g^{(2)}$ is computed and shown in (c) as a function of the integration time and fitted by Eq. (SA.41).
@@
$g^{(2)}_{\text{mes}}(\tau)=\frac{A(\tau)g^{(2)}(\tau,\Delta,\Omega)+\frac{2}{{\rm SNR}_{\text{Det}}}+\frac{1}{{\rm SNR}_{\text{Det}}^{2}}}{1+\frac{2}{{\rm SNR}_{\text{Det}}}+\frac{1}{{\rm SNR}_{\text{Det}}^{2}}},$ (SA.40)
where ${\rm SNR}_{\text{Det}}=R_{\text{sca,det}}/R_{\text{DC}}=18$ accounts for false positive photon detections due to the dark counts of the detectors. In the experiment, the dark count rate was $R_{\text{DC}}=500$ counts/s for each detector. Then, for single atoms, the photon counts are usually measured using red-detuned light in order to maintain the atom in the trap while acquiring the resonance fluorescence signal. A large negative detuning $\Delta<0$ causes an overshoot of the maximum value of $g^{(2)}$ above $2$. It is taken into account by an explicit dependence on the detuning of $g^{(2)}(\tau,\Delta,\Omega)$ which comes from the steady-state solution of the optical Bloch equations for a two-level atom. Finally, an empirical global envelop $A(\tau)=a+be^{-c\tau}$ takes into account the effect of the atomic motion in the trap Weber et al. (2006) where the parameters $a,b,c$ are extracted from a fit on microsecond timescales. After including these corrections, the correlation function is fitted from which we extract the Rabi frequency $\Omega=3.3\gamma$ and the detuning $\Delta=-3.2\gamma$. As shown on Fig. SA.4(c), modeling the atom as a two-level system including experimental imperfections is enough to explain the experimental data. This agreement justifies the use of a two-level atom model to compute the different unravelings.
@@
${\rm SNR}=\sqrt{\eta^{2}g^{(2)}R_{\text{sca,det}}t_{\text{int}}}.$ (SA.41)
Experimentally, we evaluate the signal-to-noise of the measured $g^{(2)}_{mes}$ [Fig. SA.4(c)] by computing the ratio of the average and the standard deviation in the steady-state limit [gray shaded area in Fig. SA.4(b)].
By fitting the signal-to-noise using Eq. (SA.41), we end up with $\eta=0.25\%$. First, the optical losses $\eta_{\text{loss}}=50\%$ can be improved by a better mode-matching of the emitting spatial modes and the collecting optical fibers. Second, the detector quantum efficiency $\eta_{\text{QE}}=50\%$ at a specific wavelength $\lambda=780$ nm can be increased up to $90\%$ using superconducting nanowires Natarajan et al. (2012). Finally, for an atom scattering photons isotropically, the collection efficiency $\eta_{\text{col}}$ is determined by the solid angle covered by the optical system. Based on these numbers, we deduce the experimental collection efficiency of a single lens of $1\%$. By summing all channels up, the Maltese-cross coupling scheme employing the four lenses increases in principle this coupling by a factor of 4 for an atom tightly trapped in all spatial directions. Other geometries could also be used to increase the collection efficiency, such as a parabolic mirror trap Chou et al. (2017).

### v4

@@ Abstract
Introduction.—Quantum systems interacting with Markovian environments are ubiquitous in the physical sciences. A main tool for studying their dynamics is the deterministic Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) master equation (ME) Gorini et al. (1976); Lindblad (1976). This specifies the time evolution of the density matrix $\rho(t)$ as the system experiences both coherent and incoherent processes, with the latter involving leakage of state information to the environment Breuer and Petruccione (1999); Haroche and Raimond (2006); Breuer and Petruccione (2002); Attal et al. (2006). Despite its generality and wide use, the GKSL ME does not fully describe the quantum dynamics when the environment includes measurement devices, which convert a portion of the leaked information to usable form. The temporal evolution conditioned on the measurement record $m$ defines a quantum trajectory, in the ideal case expressed as the pure state $\rho_{m}(t)=\ket{\psi_{m}(t)}\bra{\psi_{m}(t)}$ |. The unraveling into pure states can be performed whenever the continuous evolution in the GKSL ME is governed by a commutator with a non-Hermitian Hamiltonian, and the initial state is pure. The resulting nonlinear Schrödinger equation governing the evolution of normalized states in quantum-trajectory theory is a straightforward consequence of conditioning and is not considered to arise from any previously unknown inherent stochasticity Carmichael (2008). Averaging $\rho_{m}(t)$ over the measurement record solves the GKSL ME Accardi et al. (2013). The identification of the $\rho(t)$ with the ensemble average of $\rho_{m}(t)$ is an example of an unraveling of the ME into a stochastic equation for the pure state $\rho_{m}(t)$ N. Bohr and Slater (1924); Carmichael (1993); Gardiner et al. (1992); Dalibard et al. (1992); Omnès (1994); Blanchard and Jadczyk (1995); Korotkov (1999); Plenio and Knight (1998); Hegerfeldt and Sondermann (1996); Daley (2014). Evidently, different measurement schemes correspond to different unravelings and lead to different ensembles of quantum trajectories. Unravelings thus provide information that is not available from the corresponding ME.
@@
Quantities that are linear in the density matrix $\rho(t)$, such as averages of observables, are fully determined by the GKSL ME and, therefore, are independent of the choice of the unraveling dictated by a given measurement scheme. In this Letter, we develop nonlinear measures to differentiate unravelings, thus opening a way to access the physics beyond the ME. We demonstrate that evaluating an expectation value of a physical observable for a specified quantum trajectory $\rho_{m}(t)$, performing a nonlinear operation on the obtained result, and averaging the result over the measurement record, yields a quantity that allows for distinguishing different unravelings of the same GKSL ME. We focus on a paradigmatic open quantum system, the resonance fluorescence of a two-level atom, and consider unravelings corresponding to direct photodetection and to homodyne/heterodyne detection. We remark that while direct photodetection enjoys an obvious link with intensity correlations, measuring the spectrum of squeezing is inherently tied with homodyne detection, which provides a phase reference to a phase-dependent phenomenon Carmichael (2008).
@@
The unravelings.—Electron shelving Wineland and Dehmelt (1975); Dehmelt (1981) paved the way to the first observations of quantum jumps Nagourney et al. (1986); Sauter et al. (1986); Bergquist et al. (1986), followed by several atomic Basché et al. (1995); Peil and Gabrielse (1999); Gleyzes et al. (2007) and solid-state physics experiments Neumann et al. (2010); Vijay et al. (2011); Minev et al. (2019). The theoretical description of these investigations dates back to early works Cook and Kimble (1985); Kimble et al. (1986); Erber and Putterman (1985); Javanainen (1986); Schenzle et al. (1982); Cohen-Tannoudji and Dalibard (1986); Cook (1988); Grochmalicki and Lewenstein (1989a, b); Hegerfeldt and Plenio (1992, 1993); Hegerfeldt (2009) that stimulated the development of quantum trajectory theory Keys and Wehr (2020); Barchielli and Belavkin (1991); Brun (2002); Barchielli (2006); Belavkin (1990); Belavkin and Staszewski (1991); Ghirardi et al. (1990, 1986); Carmichael (1993). Two unravelings of the GKSL ME that play a fundamental role in the understanding of the quantum trajectories are: (i) Poisson unraveling, related to direct photodetection and the so called quantum Monte Carlo wave function approach Dalibard et al. (1992); Davies (1976); Srinivas and Davies (1981); Zoller et al. (1987); Dum et al. (1992); and (ii) Wiener-type unraveling (the quantum state diffusion model proposed by Gisin and Percival) Gisin and Percival (1992); Carmichael (1999), relating conditional quantum dynamics to a continuous Wiener process Durrett (2019). In the context of atomic physics experiments, the distinct unravelings correspond to different photodetection schemes Breuer and Petruccione (2002), see Fig. 1. The Poisson unraveling is relevant for the direct photodetection experiments, while the continuous Wiener process arises in homodyne and heterodyne photodection schemes Carmichael (1993); Wiseman and Milburn (1993).
@@
The disparity between the experimental setups is reflected in the different nature of quantum trajectories Carmichael (1999). The Wiener process yields a continuous evolution of the system state |$\ket{\psi(t)}$. In contrast, the acts of direct photodetection at times $t_{1}<t_{2}<\ldots<t_{n}$ collapse the conditional wavefunction. The resulting time evolution of |$\ket{\psi(t)}$ is discontinuous and the final state at $t>t_{n}$ depends, in general, on a particular sequence of emission times |$\ket{\psi(t)}=\ket{\psi_{t_{1},\ldots,t_{n}}(t)}$. By collecting photon counting records, the experimenter effectively determines the quantum trajectory of the atom. The entanglement between the electromagnetic field and the atom is the key ingredient that allows for the inference of the atom’s state based on the photodetection events Carmichael (1999); Nha and Carmichael (2004). In particular, Nha and Carmichael demonstrated that the degree of entanglement depends on how information in the environment is read Nha and Carmichael (2004).
@@
Source Master Equation and linear averages.—Our starting point is the GKSL ME of resonance fluorescence, governing the unconditional evolution of the reduced system density matrix $\rho$,
@@
The results depicted in Fig. 2 substantiate the pivotal influence of the environment when collecting records of a strongly driven two-state atom and taking a sum over a collection of them. The two principal unravelings are presented in their ability to produce an ostensibly disparate ${\rm Var}(\sigma_{z})$, while the corresponding average inversion remains unchanged. For the direct photodetection Mollow (1975); Cook (1980, 1981), corresponding to the Poisson-type unraveling of the ME, we obtain an exact expression for ${\rm Var}(\sigma_{z})$ based on the waiting-time distribution Carmichael et al. (1989); Grochmalicki and Lewenstein (1989a). For $\gamma t\gg 1$, the asymptotic expression for the variance, including first-order terms in $\gamma/\Omega$ of different frequencies, reads SM
@@
The time evolution of the QTAV is significantly altered, see the middle panels of Fig. 2, when one places a beam splitter and a local oscillator in the environment, and the fluorescent signal interferes with the latter before photodetection [Fig. 1(c)], corresponding to the heterodyne detection and exemplifying Wiener-type unraveling. The frequency doubling is also in evidence although the contrast in the oscillations is visibly suppressed. The light scattered by the two-level emitter is squeezed in the field quadrature that is in phase with the mean scattered field amplitude $\propto\langle\sigma_{-}(t)\rangle$ Walls and Zoller (1981); Mandel (1982). The bottom panel in each frame shows that the QTAV responds differently to the detection of the squeezed vs. the direction of the antisqueezed quadrature of the fluorescent field, i.e., along an axis perpendicular to the equator of the Bloch sphere where quantum fluctuations are redistributed among the quadratures.
@@
using the fact that $d\widetilde{N}d\widetilde{N}=dN$ with the rate of the Poisson process being $\gamma\left\langle\sigma_{+}\sigma_{-}\right\rangle_{t}dt$ and in the last line using $u^{j}$ as the vector corresponding to $\mathcal{L}^{\dagger}[X_{j}]$, $l$ as $\sigma_{+}\sigma_{-}$, and $v^{j}$ as $\sigma_{+}X_{j}\sigma_{-}$. This is an ordinary differential equation which however does not close since it requires higher order moments such as $\mathbb{E}x_{i}x_{j}x_{k}$. Again, using the Itô product rule we can calculate the equation for the higher order moments to obtain a system of ordinary differential equations which still do not close. We can repeat this procedure to arbitrarily high order but at some point we have to truncate. It can be shown that this truncation is linear in the moments, which allows us to solve the system using traditional methods of solving linear systems of ODEs. The Wiener case Barchielli and Belavkin (1991); Gisin and Percival (1992, 1997); Percival (1998) can be similarly approximately solved by using the Heisenberg equation Keys (2022)
@@
Since the intensity correlation of the scattered light reflects a nonexclusive probability of photocounting coincidences, the limited detector efficiency can be counterbalanced by increasing the number of photon “clicks” $N$ in the course of a long experimental run. Indeed, for the setup pictured in Fig. 1(a), we have concluded that the signal-to-noise ratio for $g^{(2)}(\gamma\tau\gg 1)$ in a window about one inverse of the coherence time scales with $\sqrt{N}$ SM . This allows the determination of ${\rm Var}(\sigma_{z})$ from single realizations as low as $10^{-3}$, which is the order of magnitude Monte-Carlo simulations indicate for $\eta\lesssim 0.05$. This order of magnitude can be increased using high numerical aperture collection systems Bruno et al. (2019); Natarajan et al. (2012) and efficient single-photon detectors Chou et al. (2017).
@@
This ability allows for an experimentally oriented test of the objective quantum state assumption via an EPR steering inequality Jones et al. (2007); Wiseman and Gambetta (2012); direct photodetection has been argued to be ‘more quantum’ than its Wiener-type counterparts Jones et al. (2007); Wiseman and Gambetta (2012); Daryanoosh and Wiseman (2014). While we have focused on two primary unravelings, the possibilities are in fact endless as different output channels open up. For example, Barchielli and Gregoratti have used a measurement-based feedback protocol to assess the non-Markovian evolution of a coherently driven and continuously monitored two-level atom. The included delay has experimental consequences, modifying the Mandel Q parameter alongside the spectrum of the emitted light Barchielli and Gregoratti (2012). In general, non-Markov open systems cannot be given a trajectory interpretation built around measured outputs, as measuring non-Markov environments interferes with the reduced system dynamics. One sub-class of trivially non-Markov open systems has been recently approached using trajectories Arranz Regidor et al. (2021). It would be interesting to extend our considerations to non-Markovian evolution case Barchielli and Gregoratti (2012). Finally, our conclusions are reflected by recent investigations of quantum many-body systems studied in the context of quantum computing and quantum simulation Preskill (2018); Altman et al. (2021); Fraxanet et al. (2023). Quantum trajectories arising due to multiple measurements of the system’s state, when analyzed by relevant (nonlinear) statistical measures such as entanglement entropy, exhibit phase transitions Skinner et al. (2019); Li et al. (2018, 2019); Chan et al. (2019); Sierant and Turkeshi (2022); Turkeshi et al. (2020); Lunt et al. (2021); Sierant et al. (2022), that are not evident in the average state Cao et al. (2019); Piccitto et al. (2022); Kolodrubetz (2023); Turkeshi et al. (2021) unless specifically tuned feedback mechanisms are employed Sierant and Turkeshi (2023a); Buchhold et al. ; Iadecola et al. (2023); Piroli et al. (2023); O’Dea et al. (2024); Ravindranath et al. (2023a, b); Sierant and Turkeshi (2023b).
@@
.7 EPR steering for system-environment entanglement
@@
$S(\rho^{c,D},\rho^{c,H})\equiv\overline{f_{1}(\rho^{c,D})}+\overline{f_{2}(\rho^{c,H})}\leq 1,$ (S.39)
@@
Figure S.3: EPR steering $S=S(\rho^{c,D},\rho^{c,H})$ value from the inequality (S.39) for: $Y=10$ and $\eta=1$ in (a), $Y=30$ and $\eta=1$ in (b), $Y=30$ and $\eta=0.8$ in (c), and $Y=30$ and $\eta=0.6$ in (d).
Figure S.3 plots the value of $S$ for different values of $Y$ and $\eta$. As noted in Wiseman and Gambetta (2012), for $S>1$ the experiment would rule out all theories of objective atomic state reduction. For unit-efficiency direct detection ($\eta=1$), the steady-state value of $S$ always exceeds unity. We note the envelope is primarily determined by the detection efficiency while the carrier frequency is set by $Y$. Figures S.3(c, d) show that the envelope drops below unity for $\eta\lesssim 0.8$, satisfying the inequality (S.39), whence making the experiment “not provably measurement dependent”. In contrast, we note that the feedback scheme adopted for direct photodetection in Wiseman and Gambetta (2012) guarantees $S>1$ for any value of $\eta$ and a perfect heterodyne detector.
.8 Experimental setup and considerations
@@
We consider real experimental data based on the setup of Bruno et al. (2019) in order to highlight the different imperfections. The experimental setup consists of a Maltese-cross coupling of single neutral atoms of 87Rb. The atom is illuminated by a laser field with Rabi frequency $\Omega$ and detuning $\Delta$. The photons are collected by high-numerical aperture lenses, as shown in the main text, and the photon arrival times are measured [Fig. S.4(a)] using single-photon avalanche photodiodes (APDs). From these measurements, the normalized second-order correlation function between the lenses L1 and L2 is computed [Fig. S.4(b)]. The average atomic scattering rate on a single detector is $R_{\text{sca,det}}=9000$ counts/s.
Figure S.4: (a) The arrival times of the photons coming from the lenses L1 and L2 are measured by avalanche photodiodes (APDs). (b) Normalized second-order correlation function for integration times of 65 (blue) or 1700 (black) seconds and fit using Eq. (S.40) from a two-level system approximation including experimental imperfections (red). The gray shaded box corresponds to the time range where the signal-to noise of $g^{(2)}$ is computed and shown in (c) as a function of the integration time and fitted by Eq. (S.41).
@@
$g^{(2)}_{\text{mes}}(\tau)=\frac{A(\tau)g^{(2)}(\tau,\Delta,\Omega)+\frac{2}{{\rm SNR}_{\text{Det}}}+\frac{1}{{\rm SNR}_{\text{Det}}^{2}}}{1+\frac{2}{{\rm SNR}_{\text{Det}}}+\frac{1}{{\rm SNR}_{\text{Det}}^{2}}},$ (S.40)
where ${\rm SNR}_{\text{Det}}=R_{\text{sca,det}}/R_{\text{DC}}=18$ accounts for false positive photon detections due to the dark counts of the detectors. In the experiment, the dark count rate was $R_{\text{DC}}=500$ counts/s for each detector. Then, for single atoms, the photon counts are usually measured using red-detuned light in order to maintain the atom in the trap while acquiring the resonance fluorescence signal. A large negative detuning $\Delta<0$ causes an overshoot of the maximum value of $g^{(2)}$ above $2$. It is taken into account by an explicit dependence on the detuning of $g^{(2)}(\tau,\Delta,\Omega)$ which comes from the steady-state solution of the optical Bloch equations for a two-level atom. Finally, an empirical global envelop $A(\tau)=a+be^{-c\tau}$ takes into account the effect of the atomic motion in the trap Weber et al. (2006) where the parameters $a,b,c$ are extracted from a fit on microsecond timescales. After including these corrections, the correlation function is fitted from which we extract the Rabi frequency $\Omega=3.3\gamma$ and the detuning $\Delta=-3.2\gamma$. As shown on Fig. S.4(c), modeling the atom as a two-level system including experimental imperfections is enough to explain the experimental data. This agreement justifies the use of a two-level atom model to compute the different unravelings.
@@
${\rm SNR}=\sqrt{\eta^{2}g^{(2)}R_{\text{sca,det}}t_{\text{int}}}.$ (S.41)
Experimentally, we evaluate the signal-to-noise of the measured $g^{(2)}_{mes}$ [Fig. S.4(c)] by computing the ratio of the average and the standard deviation in the steady-state limit [gray shaded area in Fig. S.4(b)].
By fitting the signal-to-noise using Eq. (S.41), we end up with $\eta=0.25\%$. First, the optical losses $\eta_{\text{loss}}=50\%$ can be improved by a better mode-matching of the emitting spatial modes and the collecting optical fibers. Second, the detector quantum efficiency $\eta_{\text{QE}}=50\%$ at a specific wavelength $\lambda=780$ nm can be increased up to $90\%$ using superconducting nanowires Natarajan et al. (2012). Finally, for an atom scattering photons isotropically, the collection efficiency $\eta_{\text{col}}$ is determined by the solid angle covered by the optical system. Based on these numbers, we deduce the experimental collection efficiency of a single lens of $1\%$. By summing all channels up, the Maltese-cross coupling scheme employing the four lenses increases in principle this coupling by a factor of 4 for an atom tightly trapped in all spatial directions. Other geometries could also be used to increase the collection efficiency, such as a parabolic mirror trap Chou et al. (2017).